Find the speed of a 22 kg mass with a momentum of 450 kgm/s. (Round answers to 3 significant digits as usual in all answers)
1
Expert's answer
2020-10-23T11:52:49-0400

We can find the speed of the mass from the definition of the linear momentum:


"p=mv,""v=\dfrac{p}{m}=\dfrac{450\ \dfrac{kgm}{s}}{22\ kg}=20.5\ \dfrac{m}{s}."

Answer:

"v=20.5\ \dfrac{m}{s}."